# Ali Dasdan KD Consulting Saratoga, CA, USA

### v2

@@ 4.3.2 Relaxing the Uniform Productivity Assumption
Dominance under heterogeneity. With heterogeneous productivity gains $\alpha_{i}$ varying across firms, the dominance result (Proposition 7) remains fully valid: each firm $i$ has incentive to expand given its own $\alpha_{i}>1$, regardless of others’ productivity gains. The proof is identical, replacing $\alpha$ with $\alpha_{i}$ throughout. Therefore the core mechanism driving the tragedy—individual incentive to defect regardless of what others do—persists under heterogeneous productivity shocks.
Pareto-inefficiency under heterogeneity. The Pareto-inefficiency result (Proposition 8) requires modification. With heterogeneous $\alpha_{i}$, workers remain strictly better off under $s^{C}$, but the outcome for firms depends on the distribution of productivity gains:
$U_{F,i}(s^{C})-U_{F,i}(s^{D})$ $=\left(\frac{Q_{i}}{\sum_{j}Q_{j}}-\frac{\alpha_{i}Q_{i}}{\sum_{j}\alpha_{j}Q_{j}}\right)+\left(-c\cdot\frac{H_{i}}{\alpha_{i}}\right)-\left(-c\cdot H_{i}\right)$
$=\underbrace{\pi\cdot Q_{i}\left(\frac{\sum_{j}(\alpha_{j}-\alpha_{i})Q_{j}}{(\sum_{j}Q_{j})\cdot(\sum_{j}\alpha_{j}Q_{j})}\right)}_{\text{Market Share Change}}+\underbrace{c\cdot\frac{(\alpha_{i}-1)H_{i}}{\alpha_{i}}}_{\text{Labor Cost Savings}}$
Firms with higher productivity gains capture larger market shares under universal defection. Whether universal cooperation Pareto-dominates then depends on the distribution of productivity gains. If productivity differences are modest, cooperation can still dominate. If productivity differences are extreme, some high-productivity firms might prefer universal defection as it allows market share capture from lower-productivity rivals.
Critically, heterogeneity makes coordination more difficult, not easier. Firms with productivity advantages face even stronger temptation to defect and capture market share from lower-productivity rivals. This reinforces rather than undermines our structural analysis, as the tragedy becomes more severe when productivity gains are unequally distributed.
@@ 4.3.3 Generality and Robustness
Beyond binary choices. The dominance property is not an artifact of the binary choice model. It extends to continuous strategy spaces where firms choose specific hours $h_{i}\in[0,H_{max}]$. In such settings, the marginal incentive to increase output (to capture market share) persists at every level of $h_{i}$ provided the competitive elasticity is sufficiently high.
Risk aversion strengthens the tragedy. Since the utility loss from losing market share (existential risk) outweighs the utility gain from leisure, risk-averse firms face even stronger pressures to defect than risk-neutral ones.
Generality across market structures. The tragedy holds for:
@@
  • Heterogeneous productivity gains $\alpha_{i}$ across firms
@@ 4.4 Empirical Validation: The European Case
The European experience provides crucial empirical validation as a best-case scenario. If our structural analysis is correct, even under maximally favorable conditions for coordination—wealthy democracies with strong institutions—we should observe:
  1. 1.
    Partial override: coordination achieves some welfare improvements but not full Keynesian leisure
  2. 2.
    High costs: coordination requires substantial institutional apparatus
  3. 3.
    Erosion over time: competitive pressure gradually undermines coordination
The evidence strongly supports these predictions.
Favorable Conditions.
Europe represents the most favorable case for coordination:
  • Wealthy democracies with strong institutions
  • Supranational governance (European Union) providing coordination capacity
  • Cultural consensus supporting work-life balance
  • Strong labor unions with political influence
  • Legal frameworks mandating maximum hours and minimum vacation
If coordination is possible anywhere, it should succeed in Europe.
Partial Override.
European workers have achieved shorter hours than American workers, e.g., about 1,350 hours in Germany vs. about 1,800 hours in the US (OECD, 2025). Workers in many developing economies face even longer hours with weaker protections, demonstrating that the coordination failure intensifies under less favorable institutional conditions.
@@
This represents real achievement: European coordination reduced hours by approximately 450 compared to the United States. However, even German workers work far more than Keynes’s predicted 15 hours weekly (780 hours annually). The coordination is partial, not complete.
Moreover, productivity-welfare decoupling manifests differently across institutional contexts. While European workers gained some leisure relative to their US counterparts, Europe also exhibits:
  • Declining labor share of income since 1980, similar to the US trend
  • Slower wage growth relative to productivity in many sectors
  • Persistent pressure on social benefits and employment protections
  • Higher unemployment rates in several major economies
@@
The tragedy is not escaped but redistributed: European workers gained some leisure at the cost of wage growth and employment, while US workers experienced stagnant wages and benefits despite longer hours. Both regions show the fundamental pattern our framework predicts: productivity-welfare decoupling driven by competitive pressure, manifesting through different dimensions (hours vs. wages vs. benefits vs. job security) depending on institutional context.
High Costs.
This coordination came with substantial costs:
@@
Economic costs: European GDP growth has lagged the United States annually over recent decades; for example, data indicates a significant divergence in aggregate growth trajectories between EU and the USA since 2008 (World Bank, 2024). While this divergence is driven by multiple factors, shorter working hours impose a mechanical ceiling on total output.
@@
Regulatory costs: Maintaining shorter hours requires extensive regulation. France’s 35-hour workweek involves complex rules about overtime, exceptions, and compliance monitoring (L&E Global, 2024). Germany’s system involves workplace councils, sector-level bargaining, and intricate labor law.
@@
Competitive pressure: European firms face disadvantages competing against firms from regions with longer working hours. This shows up in market share, innovation metrics, and firm valuations. The competitive pressure remains even under coordination.
Erosion Over Time.
Most tellingly, the coordination shows signs of erosion:
Enforcement weakening: France’s 35-hour law contains numerous exceptions and workarounds. Actual hours often exceed nominal limits through overtime, managerial exemptions, and definitional ambiguities (L&E Global, 2024).
Pressure for reform: Germany’s Hartz reforms increased flexibility and effectively lengthened working hours for some sectors (IAB-Forum, 2023).
Competitive concerns: European policy discussions increasingly focus on competitiveness relative to the United States and China, with pressure to loosen labor regulations to boost growth.
@@
Hours stabilization, not decline: European working hours have largely stabilized since the 1980s despite continued productivity growth (Huberman and Minns, 2007). The decline in hours observed from 1870-1980 has stalled, suggesting coordination has reached its limits.
Interpretation.
The European case validates our framework in three ways:
First, even under maximally favorable conditions, coordination is partial. European workers still work far more than technological capacity allows. The tragedy is attenuated but not resolved.
Second, coordination is costly. Substantial institutional apparatus is required, and even this apparatus faces continuous pressure and erosion. The coordination is fragile, not robust.
Third, competitive pressure persists and manifests across multiple dimensions. European firms face disadvantages that generate political pressure to loosen regulations. Even where hours were reduced, other dimensions of worker welfare (wages, employment, benefits) came under pressure. The underlying tragic structure remains even when temporarily overridden by institutions.
This shows that the productivity tragedy is structural, not merely cultural or policy-related. Even when culture supports leisure, institutions enable coordination, and wealth permits reduced hours, the competitive structure creates barriers that limit welfare improvements and erode over time.
4.5 Why Standard Solutions Face Structural Barriers
Our structural analysis explains why proposed solutions face barriers:
Solution 1: Individual Firm Action
@@ 4.5 Why Standard Solutions Face Structural Barriers
Problem: Violates dominance condition. Any firm reducing hours unilaterally suffers competitive disadvantage. Competitors capture market share, potentially forcing the cooperative firm out of business. First-mover disadvantage is severe.
Solution 2: Industry-Level Agreements
Problem: Violates enforcement condition. Agreements lack binding force absent government enforcement. Each firm has incentive to defect secretly. Even with monitoring, defection detection may come too late to prevent competitive harm.
Solution 3: National Regulation
Problem: Faces international competition. Capital mobility enables firms to relocate to jurisdictions with longer hours or lower labor costs. This race-to-the-bottom dynamic affects both wealthy and developing nations, with the latter often experiencing more intense pressure to attract investment through weaker labor protections.
Solution 4: International Agreements
Problem: Violates enforcement condition. International labor standards (ILO conventions) lack enforcement mechanisms. Compliance is voluntary. Countries competing for investment have incentives to defect through weak enforcement or exceptions.
Solution 5: Cultural Change
Problem: Insufficient against structural pressure. Even if workers prefer leisure, competitive pressure compels long hours. Firms facing bankruptcy cannot afford to honor worker preferences. Cultural change may be necessary but is insufficient absent structural change.
Solution 6: Technological Unemployment
Problem: Historical evidence suggests employment adjusts rather than hours reducing. When technology displaces workers from one sector, they typically move to other sectors rather than working less (Autor, 2015). Unemployment may increase but this differs from coordinated work reduction.
These solutions all fail because they don’t address the underlying structure. The tragedy persists because the five conditions remain satisfied. Only interventions that break at least one condition could resolve the tragedy.
4.6 What Would Work
Based on our framework, effective interventions must break at least one of the five conditions:
Break Condition 1 (N-Player): Reduce the number of independent actors. This is difficult—markets inherently involve multiple firms, and consolidation faces antitrust constraints. Complete monopoly might enable work reduction but at costs to competition and innovation.
Break Condition 2 (Externalities): Eliminate the externality where one firm’s expansion harms others. This seems impossible since competitive pressure is inherent to markets. Work reduction by one firm creates market share opportunities for others by definition.
Break Condition 3 (Dominance): Reduce existential stakes making defection dominant. This is most promising. If firms didn’t face bankruptcy risk from work reduction, coordination becomes feasible. Mechanisms:
  • Universal Basic Income reduces worker pressure for long hours
  • Stronger social safety nets reduce survival stakes
  • Reduced international competition lessens competitive pressure
  • Market structures that reward restraint (currently absent)
Break Condition 4 (Pareto-Inefficiency): Accept the equilibrium rather than seeking coordination. Some economists argue long hours reflect preferences rather than coordination failure (Pencavel, 2018). Our framework suggests this is incorrect as the Pareto-domination is real, and workers would benefit from coordination.
Break Condition 5 (Enforcement): Solve the enforcement problem through effective global labor standards with monitoring and sanctions. This requires:
  • International agreement on maximum hours
  • Verification mechanisms for compliance
  • Sanctions for defection sufficient to deter
  • Sufficient participation to prevent race-to-bottom
This is theoretically possible but practically difficult given international anarchy and sovereignty concerns.
The productivity tragedy can be solved, but only through interventions that address its structural foundations. Incremental reforms that leave the five conditions intact will achieve at most partial, costly, eroding coordination as observed in Europe.

### v3

@@ 4.3.2 Relaxing the Uniform Productivity Assumption
Dominance Under Heterogeneity:
With heterogeneous productivity gains $\alpha_{i}$ varying across firms, the dominance result (Proposition 7) remains fully valid: each firm $i$ has incentive to expand given its own $\alpha_{i}>1$, regardless of others’ productivity gains. The proof is identical, replacing $\alpha$ with $\alpha_{i}$ throughout. Therefore the core mechanism driving the tragedy—individual incentive to defect regardless of what others do—persists under heterogeneous productivity shocks.
Pareto-inefficiency Under Heterogeneity:
The Pareto-inefficiency result (Proposition 8) assumes comparable productivity gains. With heterogeneous $\alpha_{i}$, workers remain strictly better off under $s^{C}$, but the outcome for firms depends on the distribution of productivity gains.
When heterogeneity is extreme, high-productivity firms may prefer universal defection to capture market share from rivals. However, such extreme heterogeneity leads to rapid market consolidation through the exit of low-productivity firms (Melitz, 2003). The surviving firms then face the tragedy among themselves: once similar-productivity firms remain in the market, Pareto-inefficiency re-emerges among the competitive set. Thus, extreme heterogeneity does not eliminate the tragedy but rather localizes it to the relevant competitive cohort.
For the empirically common case of moderate heterogeneity within industries, Pareto-inefficiency holds: the productivity distribution of surviving firms exhibits substantial but not extreme dispersion (Syverson, 2011), and all firms benefit from coordinated restraint relative to the racing equilibrium. The dominance property (Condition 3) ensures that even when C4 holds weakly, coordination remains structurally difficult because every firm prefers unilateral defection.
@@ 4.3.3 Generality and Robustness
Beyond Binary Choices:
The dominance property is not an artifact of the binary choice model. It extends to continuous strategy spaces where firms choose specific hours $h_{i}\in[0,H_{max}]$. In such settings, the marginal incentive to increase output (to capture market share) persists at every level of $h_{i}$ provided the competitive elasticity is sufficiently high.
Risk Aversion Strengthens the Tragedy:
Since the utility loss from losing market share (existential risk) outweighs the utility gain from leisure, risk-averse firms face even stronger pressures to defect than risk-neutral ones.
Generality Across Market Structures:
The tragedy holds for:
@@
  • Heterogeneous productivity gains $\alpha_{i}$ (within the surviving competitive cohort)
@@ 4.4 Empirical Validation: The European Case
The European experience provides crucial empirical validation. If our structural analysis is correct, even under maximally favorable institutional conditions, we should observe: (1) partial mitigation achieving some welfare improvements but not full Keynesian leisure, (2) high costs requiring substantial institutional apparatus, and (3) erosion over time as competitive pressure gradually undermines coordination.
Favorable Conditions:
Europe represents the most favorable case for coordination: wealthy democracies with strong institutions, supranational governance (European Union) providing coordination capacity, cultural consensus supporting work-life balance, strong labor unions with political influence, and legal frameworks mandating maximum hours and minimum vacation. If coordination is possible anywhere, it should succeed in Europe.
Partial Mitigation:
@@
European workers have achieved shorter hours than American workers—approximately 1,350 hours annually in Germany versus 1,800 in the US (OECD, 2025). This represents real achievement: a reduction of approximately 450 hours compared to the United States. However, even German workers work far more than Keynes’s predicted 780 hours annually. The coordination is partial, not complete.
@@
Moreover, productivity-welfare decoupling manifests differently across institutional contexts. While European workers gained leisure relative to US counterparts, Europe also exhibits declining labor share of income since 1980, slower wage growth relative to productivity, persistent pressure on social benefits, and higher unemployment in several major economies. The tragedy is not escaped but redistributed: European workers gained leisure at the cost of wage growth and employment security, while US workers experienced stagnant wages despite longer hours. Both regions show productivity-welfare decoupling driven by competitive pressure, manifesting through whichever dimension institutional constraints leave exposed.
High Costs:
This coordination requires substantial costs.
@@
Economic costs: European GDP growth has lagged the United States over recent decades, with significant divergence since 2008 (World Bank, 2024). While this divergence is driven by multiple factors, shorter working hours impose a mechanical ceiling on total output.
@@
Regulatory costs: Maintaining shorter hours requires extensive regulation: France’s 35-hour workweek involves complex rules about overtime, exceptions, and compliance monitoring (L&E Global, 2024), while Germany’s system requires workplace councils, sector-level bargaining, and intricate labor law.
@@
Competitive pressure: European firms face competitive disadvantages against firms from regions with longer working hours, generating continuous political pressure to loosen regulations.
Erosion Over Time:
@@
Most tellingly, coordination shows signs of erosion. France’s 35-hour law contains numerous exceptions and workarounds; actual hours often exceed nominal limits (L&E Global, 2024). Germany’s Hartz reforms increased flexibility and effectively lengthened working hours in some sectors (IAB-Forum, 2023). European working hours have largely stabilized since the 1980s despite continued productivity growth (Huberman and Minns, 2007); the decline observed from 1870-1980 has stalled, suggesting coordination has reached its limits under persistent competitive pressure.
Interpretation:
The European case demonstrates that institutional intervention can achieve partial, costly mitigation—not elimination of the tragedy. Institutions shift the equilibrium toward Pareto-optimality but cannot reach it absent structural change to competitive conditions themselves. Even under maximally favorable conditions, coordination remains partial (hours far exceed technological capacity), costly (requiring extensive institutional apparatus), and fragile (showing continuous erosion). The underlying tragic structure persists even when temporarily overridden.
4.5 Why Standard Solutions Face Structural Barriers?
Our framework explains why commonly proposed solutions to the productivity-welfare decoupling face structural barriers. Each proposed solution fails to address one or more of the five necessary conditions.
Solution 1 — Individual Firm Action:
Encourage firms to voluntarily reduce working hours or increase wages proportional to productivity gains.
@@ 4.5 Why Standard Solutions Face Structural Barriers?
Problem: Violates dominance condition (C3). Any firm reducing hours unilaterally suffers competitive disadvantage while competitors capture its market share. The logic of collective action (Olson, 1965) explains why voluntary restraint by individual firms cannot succeed when defection yields higher payoffs regardless of competitors’ strategies. First-mover disadvantage is severe in competitive markets.
Solution 2 — Industry-Level Voluntary Agreements:
Create voluntary agreements among firms to coordinate on reduced hours or higher wages.
Problem: Violates enforcement condition (C5). As Stigler (1964) demonstrated for cartel stability, voluntary agreements among competitors are inherently unstable because each firm has a dominant incentive to defect by secretly expanding hours to capture market share while others comply. While reputational mechanisms or industry associations might provide weak enforcement, these cannot overcome the profit advantages from defection. Without government enforcement imposing credible sanctions, such agreements inevitably collapse.
Solution 3 — National Regulation:
Implement national labor standards requiring reduced working hours or minimum wage increases tied to productivity.
Problem: Faces international competition and capital mobility. Böckerman and Maliranta (2012) provide micro-level evidence from Finland: despite strong labor market institutions and centralized wage bargaining, exposure to international trade systematically drove exits of high-labor-share plants, reducing the aggregate labor share by reallocating production toward more competitive establishments. This validates the race-to-the-bottom dynamic (Rodrik, 1997), demonstrating that even wealthy democracies with strong institutions face structural pressure from international competition. National regulation without global coordination merely shifts production to less regulated jurisdictions.
Solution 4 — International Labor Agreements:
Establish international treaties or agreements on labor standards, including maximum working hours and minimum wages tied to productivity.
Problem: Violates enforcement condition (C5). Flanagan (2003) demonstrates that international labor standards face severe enforcement barriers: the International Labour Organization lacks sanctioning power, compliance is voluntary, and nations face strong incentives to free-ride by endorsing standards publicly while maintaining lax enforcement to attract capital investment. As Ostrom (1990) established, effective governance requires graduated sanctions and low-cost monitoring, both absent at the international level due to sovereignty concerns and international anarchy.
Solution 5 — Technological Unemployment:
Assume automation will eventually force shorter working weeks as human labor becomes unnecessary.
Problem: Misidentifies the mechanism. As Autor (2015) documents, despite dramatic technological progress, employment has remained robust because technology creates complementary tasks even as it automates existing ones. More fundamentally, automation increases productivity, which intensifies rather than resolves the coordination failure. The dominance condition (C3) ensures that firms use productivity gains to expand market share rather than reduce hours, unless coordination mechanisms force otherwise. Historical evidence supports this: productivity increased eightfold since Keynes’s prediction, yet working hours declined far less than proportionally.
Solution 6 — Cultural Change:
Encourage societal preference shifts toward valuing leisure over consumption.
Problem: Insufficient against structural pressure. Even if workers prefer leisure, competitive dynamics compel long hours. Schor (1992) documents that despite expressed preferences for more leisure time, American workers face structural pressures that override these preferences. The dominance condition (C3) ensures that cultural preferences alone cannot overcome the structural necessity of maintaining competitive position when others work longer hours. Preferences must align with incentives, which the competitive structure prevents.
These solutions face structural barriers because they fail to address the complete set of conditions. Partial solutions that leave some conditions unsatisfied achieve at most temporary, costly, and eroding coordination, exactly as observed in Europe.
4.6 What Would Work?
Our framework identifies interventions that could resolve the productivity tragedy by breaking one or more of the five necessary conditions. We emphasize that these interventions are structural requirements, not policy recommendations, to clarify the magnitude of barriers any solution must overcome.
Break Condition 1 (N-Player Structure):
Consolidate decision-makers into smaller groups, making coordination feasible through reduced free-rider problems. As Olson (1965) established, smaller groups ($n$) dramatically lower coordination costs.
However, achieving this through industrial consolidation faces severe barriers. Antitrust law in most jurisdictions prohibits anticompetitive coordination, and for good reason: consolidation that enables coordination on labor practices would simultaneously enable price-fixing and output restriction, trading one market failure (coordination tragedy) for another (monopoly power) that harms consumer welfare. Alternatively, international consolidation of labor regulation authority could reduce the number of competing jurisdictions, but this faces political barriers from sovereignty concerns and competing national interests.
Break Condition 2 (Binary Choice with Externalities):
Eliminate the competitive externality where one firm’s expansion harms competitors’ market positions.
However, this is structurally impossible since competitive pressure is inherent to market economies. As Schumpeter (1942) argued, competition through creative destruction—where successful firms displace less efficient ones—is capitalism’s fundamental mechanism for driving productivity growth. Removing this externality would eliminate the competitive process itself. The externality is reciprocal: protecting compliant firms from market share loss necessitates restricting expanding firms, which contradicts the productivity-enhancing function of market competition.
Break Condition 3 (Dominance Property):
Reduce the existential stakes of restraint so that cooperation doesn’t threaten firm or worker survival.
For workers, mechanisms include Universal Basic Income (UBI), stronger unemployment insurance, or other social safety nets that provide unconditional material security. As Van Parijs (1995) argues, guaranteeing an unconditional material baseline grants workers “real freedom” to refuse exploitative employment terms, breaking the dominance of survival pressure that compels acceptance of maximum-intensity work arrangements.
For firms, mechanisms would include subsidies, tax incentives, or market structures that reward restraint rather than penalizing it; for example, government contracts preferentially awarded to firms with reduced hours, or subsidies compensating for reduced output during hour reductions.
However, both face substantial implementation barriers. For worker-side mechanisms, fiscally adequate UBI levels require substantial tax increases or redistribution that face political opposition across diverse political systems (Martinelli, 2017). Even where safety nets exist, they may be insufficient to overcome competitive pressure when rivals maintain longer hours; workers with basic income still face relative income competition and status concerns that perpetuate long hours (Frank, 1999).
For firm-side mechanisms, subsidies reducing bankruptcy risk from reduced output face severe moral hazard concerns: firms might claim hour reductions while maintaining actual hours, or exploit subsidies without genuine coordination. Additionally, such programs would require ongoing substantial public expense and create competitive distortions favoring subsidized over non-subsidized firms. Most fundamentally, any mechanism rewarding restraint must overcome the problem that market structures inherently reward expansion—changing this requires fundamentally restructuring competitive dynamics, which returns to the impossibility of breaking Condition 2 (eliminating competitive externalities).
Break Condition 4 (Pareto-Inefficiency):
Accept the equilibrium as efficient rather than seeking coordination. Some economists, notably Prescott (2004), argue that transatlantic differences in working hours reflect rational responses to marginal tax rate differences rather than coordination failures, with Americans optimally choosing longer hours given higher after-tax wages.
However, as Pencavel (2018) demonstrates, observed hours often do not reflect worker preferences; instead, institutional and market constraints significantly determine working time, and workers cannot simply choose their preferred hours. Additionally, the tax-incentive explanation cannot account for the secular decline in labor share within countries over time, including during periods of tax stability. The micro-level restructuring documented by Böckerman and Maliranta (2012)—where high-labor-share plants systematically exit under competitive pressure—provides direct evidence of coordination failure operating alongside any tax or preference effects. If workers genuinely prefer current hours and the equilibrium is Pareto-efficient, no coordination problem exists. Our framework and the empirical evidence suggest otherwise.
Break Condition 5 (Enforcement Difficulty):
Solve the enforcement problem through effective global labor standards with comprehensive monitoring and binding sanctions. This requires international agreement on maximum hours and minimum wage-productivity linkages, verification mechanisms for compliance across jurisdictions, sanctions sufficient to deter defection, and sufficient participation to prevent race-to-bottom through non-compliant jurisdictions.
However, effective global enforcement faces formidable barriers. As Ostrom (1990) identified, successful commons governance requires graduated sanctions and low-cost monitoring—both absent in global labor markets due to international anarchy, sovereignty concerns, and capital mobility that enables regulatory arbitrage. The International Labour Organization establishes standards but lacks enforcement power; achieving binding global labor standards would require unprecedented coordination among competing nations, effectively solving a higher-order coordination problem to address the first-order productivity tragedy.
Comprehensive Solution:
The productivity tragedy can be solved, but only through interventions that address its structural foundations. Given the difficulty of breaking any single condition completely, the most feasible approach likely combines partial interventions across multiple conditions: strengthening social safety nets (C3), enhancing coordinated regional labor standards (C5) among economically integrated blocs, and accepting modest efficiency costs from reduced competitive pressure. However, as European experience demonstrates, even such comprehensive efforts face continuous erosion from global competitive pressure. Incremental reforms that leave the five conditions substantially intact will achieve at most partial, costly, and eroding coordination.